OpenStudy (anonymous):

Find the zeros (roots) of the quadratic equation. x2+8x−20=0 A. - 5, - 15 B. - 20, 10 C. - 10, 2

OpenStudy (anonymous):

split as x2+10x-2x-20=0 x(x+10)-2(x+10)=0 (x-2)(x+10)=0

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

x=2 and -10